OpenStudy (anonymous):

Find the three arithmetic means between 10 and 18.

OpenStudy (anonymous):

First AM=(10+18)/2 = (28)/2 =14 Second AM=(10+14)/2 = (24)/2 =12 Third AM=(10+12)/2 = (22)/2 =11 Hence 11, 12, 14 are the three AM between 10 and 18 and are written as 10,11, 12, 14, 18